OpenStudy (anonymous):

Find a general form equation for the line through the pair of points (4,2) and (8,-2)

OpenStudy (gorv):

we need to apply poin slope form

OpenStudy (gorv):

\[y-y1=\frac{ y2-y1 }{ x2-x1 }*(x-x1)\]

OpenStudy (gorv):

can u do now ??

OpenStudy (gorv):

out of two point consider one as (x1,y1) and other one (x2,y2)

OpenStudy (wolf1728):

(4,2) (8,-2) slope = (-2 -2) / (8 -4) slope = -4/-4 = -1 A linear equation is written y = mx + b 'm' is the slope and so at this point we can say the equation is y = -x + b Let's rewrite this as b = y +x Let's take one of the original points (4, 2) and plug it in and we get b = y + x b = 6 Therefore the complete equation is y = -x +6 (Here's a calculator to check that answer:) http://www.1728.org/distance.htm

OpenStudy (wolf1728):

oops I made a mistake way at the top slope = -4 / 4 = -1 Everything else is fine